# Myers–Steenrod theorems for metric and singular Riemannian foliations

### v1

@@ 2.1. Singular Riemannian foliations
    For each point $x\in M$, there exists a smooth family of vector fields spanning $T_{x}L_{x}$, the tangent space to $L_{x}$ at $x$.
@@
We call any leaf of maximal dimension a regular leaf; leaves that are not regular are called singular. Given any Riemannian manifold $M$, the foliation consisting of one leaf =$\mathcal{F}=\{M\}$ and the foliation where each leaf consists of just a point =$\mathcal{F}=\{\{x\}\mid x\in M\}$ are trivial examples of singular Riemannian foliations. Other examples of singular Riemannian foliations are given by the partition of a complete Riemannian manifold $M$ by the orbits of an isometric action of a compact Lie group. The partition of $M$ into the fibers of a Riemannian submersion $f\colon M\to N$ yields a further example of a singular Riemannian foliation. Note that there are infinitely many examples of singular Riemannian foliations which not given by group actions nor from Riemannian submersions [16, 38].
@@
Let $(M,\mathcal{F})$ be a singular Riemannian foliation. Then we have a singular distribution $H\subset M$, called the horizontal distribution, given by setting $H_{x}=\nu_{x}L_{x}$, the normal tangent space to the leaf $L_{x}$ at $x$. The codimension of the foliation, denoted by $\operatorname{codim}(\mathcal{F})$, is the codimension of any regular leaf in $M$. If all the leaves of the foliation are closed submanifolds, we say that the foliation is closed. The leaf space of the foliation is the set of equivalence classes $M^{\ast}=M/\mathcal{F}$ equipped with the quotient topology. We have a natural projection map $\pi\colon M\to M^{\ast}$ which is continuous with respect to the quotient topology. Given a subset $A\subset M$, we let $A^{\ast}=\pi(A)$.
@@
Let $(M,\mathcal{F})$ be a closed singular Riemannian foliation. Fix $x\in M$ and consider the normal tangent space $\nu_{x}L_{x}$ to the leaf $L_{x}\subset M$. Next, for $\varepsilon>0$ sufficiently small, define $\nu_{x}^{\varepsilon}L_{x}=\exp_{x}((\nu_{x}L_{x})\cap B_{\varepsilon}(0))$, where $B_{\varepsilon}(0)$ is the closed ball of radius $\varepsilon$ in $T_{x}M$. Set $S_{x}=\exp_{p}(\nu_{x}^{\varepsilon}L_{x})$. The intersection of the leaves of $\mathcal{F}$ with $S_{x}$ induces a foliation $\mathcal{F}|_{S_{x}}$ on $S_{x}$ whose leaves are the connected components of the intersection between the leaves of $\mathcal{F}$ and $S_{x}$. Although $\mathcal{F}|{S_{x}}$ may not be a singular Riemannian foliation with respect to the induced metric of $M$ on $S_{x}$ (the leaves of $\mathcal{F}|{S_{x}}$ may not be equidistant with respect to the induced metric), the pull-back foliation $\mathcal{F}^{x}=\exp_{x}^{\ast}(\mathcal{F}|_{S_{x}})$ is a singular Riemannian foliation on $\nu_{x}^{\varepsilon}L_{x}$ equipped with the Euclidean metric (see [35, Proposition 6.5]). The foliation $(\nu_{x}^{\varepsilon}L_{x},\mathcal{F}^{x})$ is called the infinitesimal foliation at $x$.
@@
The infinitesimal foliation $(\nu_{x}^{\varepsilon}L_{x},\mathcal{F}^{x})$ is invariant under homotheties fixing the origin (see [35, Lemma 6.2]). Furthermore, the origin $\{0\}\subset\nu_{x}^{\varepsilon}L_{x}$ is a leaf of the infinitesimal foliation. Since the leaves of $\mathcal{F}^{x}$ are equidistant, the origin being a leaf implies that any leaf of $\mathcal{F}^{x}$ is at a constant distance from $\{0\}$. Therefore, each leaf of the infinitesimal foliation is contained in a round sphere centered at the origin. Hence, we may consider the infinitesimal foliation restricted to the unit normal sphere of $\nu_{x}L_{x}$, denoted by $S^{\bot}_{x}$, resulting in a foliated round sphere $(S^{\bot}_{x},\mathcal{F}^{x})$ with respect to the standard round metric of $S^{\bot}_{x}$. This foliation is also called the infinitesimal foliation. Henceforth, when referring to the ”infinitesimal foliation”, we will mean $(S^{\bot}_{x},\mathcal{F}^{x})$. Note that there is no loss of generality in doing so since $(\nu_{x}L_{x},\mathcal{F}^{x})$ is invariant under homothetic transformations and thus one may recover it from $(S^{\bot}_{x},\mathcal{F}^{x})$.
@@
We denote by $\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})$ the group of foliated isometries of the infinitesimal foliation, i.e., the isometries which preserve the infinitesimal foliation. For each loop $\gamma$ at $x$, the map $G_{\gamma}\colon\nu_{x}L\to\nu_{x}L$ given by $G_{\gamma}(v)=G(1,v)$ is a foliated linear isometry (see [34, Corollary 1.5]). Therefore, we have a group homomorphism $\rho\colon\Omega(L,x)\to\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})$ from the loop space of $L_{x}$ at $x$ to the foliated isometries of the infinitesimal foliation by setting $\rho(\gamma)=G_{\gamma}$.
@@
An isometry in $\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})$ may map a leaf to a different leaf. By $\mathrm{O}(\mathcal{F}^{p})$ we denote the foliated isometries preserving the foliation, i.e., isometries $f\in\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})$ such that for any leaf $\mathcal{L}$ of $(S^{\perp}_{x},\mathcal{F}^{x})$, we have $f(\mathcal{L})\subset\mathcal{L}$. The natural action of $\mathrm{O}(S^{\perp}_{p},\mathcal{F}^{x})$ on the quotient $S^{\perp}_{x}/\mathcal{F}^{x}$ has kernel $\mathrm{O}(\mathcal{F}^{x})$. Using the fact that if two loops $\gamma_{1}$ and $\gamma_{2}$ based at $x$ are homotopic, then $G_{\gamma_{1}}^{-1}\circ G_{\gamma_{2}}$ is in the kernel of the action of $\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})$ on $S^{\perp}_{x}/\mathcal{F}^{x}$, one may show that there is a well-defined group homomorphism
$\rho\colon\pi_{1}(L,x)\to\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})/\mathrm{O}(\mathcal{F}^{x}),$
@@
given by $\rho[\gamma]=[G_{\gamma}]$ (see, for example, [10, Lemma 2.4 and Proposition 2.5]). We define the holonomy of the leaf $L$ as the image $\Gamma_{L}<\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})/\mathrm{O}(\mathcal{F}^{p})$ of $\pi_{1}(L,p)$ under the homomorphism $\rho$. When we consider the holonomy of a leaf $L_{x}$ trough a point $x\in M$, we denote it by $\Gamma_{x}$. A regular leaf $L$ is a principal leaf if its holonomy group is trivial, and exceptional otherwise. The set $M_{\mathrm{prin}}\subset M$ consisting of the union of prinicpal leaves is an open, connected, and dense subset of $M$ (see [10, Theorem 3.12]).
2.2. Alexandrov spaces
An Alexandrov space $(X,d)$ with curvature bounded below by $k\in\mathbb{R}$ is a complete length space of finite Hausdorff dimension with curvature bounded below in the triangle comparison sense. Namely, for each $x\in X$, there is an open neighborhood $U\subset X$ of $x$, such that, for each geodesic triangle $\triangle$ contained in $U$, there exists a geodesic triangle $\widetilde{\triangle}$ in $M^{2}_{k}$, the $2$-dimensional model space of constant sectional curvature $k$, with edges having the same lengths as the edges of $\triangle$ satisfying the following condition: If $y$ is a vertex of $\triangle$, $\widetilde{y}$ is the corresponding vertex in $\widetilde{\triangle}$, and $w$ is any point in the opposing edge in $\triangle$ with corresponding point $\widetilde{w}$ in the opposing edge in $\widetilde{\triangle}$, then
$d(y,w)\geqslant d_{M^{2}_{k}}(\widetilde{y},\widetilde{w}).$
As is customary, we will abbreviate curvature bounded below by $k$ by writing $\operatorname{curv}\geqslant k$. A complete Riemannian manifold $M$ with a uniform lower bound for the sectional curvature is an example of an Alexandrov space.
@@ 2.2. Alexandrov spaces
$\angle(c_{1},c_{2})=\lim_{s,t\to 0}\measuredangle(c_{1}(s),x,c_{2}(x)),$
provided the limit exists. When $X$ is an Alexandrov space and the curves $c_{1}$ and $c_{2}$ are geodesics, the angle $\angle(c_{1},c_{2})$ exists (see [8, Proposition 4.3.2]).
Given two geodesics $c_{1}\colon[0,1]\to X$ and $c_{2}\colon[0,1]\to X$ in an Alexandrov space $X$ with common start point $x\in X$, we say that $c_{1}$ is equivalent to $c_{2}$ if $\angle(c_{1},c_{2})=0$. Let $\widetilde{\Sigma}_{x}(X)$ the set of equivalence classes of geodesics starting at $x$. We define a metric on this set by setting the distance between two classes to be the angle formed between any two representatives of each class. The space of directions $\Sigma_{x}(X)$ of $X$ at $x$ is the metric completion of $\widetilde{\Sigma}_{x}(X)$. By Corollaries 7.10 and 7.11 in [9], the metric space $\Sigma_{x}(X)$ is an Alexandrov space of $\operatorname{curv}\geqslant 1$.
By Theorem 10.4.1 in [8], any Alexandrov space $\Sigma$ with $\operatorname{curv}\geqslant 1$ has $\operatorname{diam}(\Sigma)\leq\pi$. Let $(X,d)$ be a metric space with $\operatorname{diam}(X)\leq\pi$. The euclidean cone over $X$, denoted by $CX$, is the set $X\times[0,\infty)/(x,0)\sim(y,0)$ equipped with the metric
@@
Given an Alexandrov space $X$ and a point $x\in X$, we define the tangent cone of $X$ at $x$ as $T_{x}X=C\Sigma_{x}(X)$. We denote the vertex of $T_{x}X$ by 00.
@@
The leaf space $M^{\ast}$ of a singular Riemannian foliation $(M,\mathcal{F})$ on a complete Riemannian manifold with closed leaves inherits a complete metric $d^{\ast}$ from $M$, known as the leaf-projection metric. For $x^{\ast},y^{\ast}\in M^{\ast}$, the distance $d^{\ast}(x^{\ast},y^{\ast})$ is defined as
$d^{\ast}(x^{\ast},y^{\ast})=d(L_{x},L_{y}),$
the distance between the leaves $L_{x}$ and $L_{y}$ considered as subsets of $M$. Equipped with the metric $d^{\ast}$, the leaf space $M^{\ast}$ has curvature locally bounded below in the triangle comparison sense discussed above. Specifically, if $U\subset M$ is an open neighborhood with sectional curvature bounded below by $k_{U}\in\mathbb{R}$, then the projection $U^{\ast}\subset M^{\ast}$ has curvature bounded below by $k_{U}$ (see [31]). The Hausdorff dimension of $M^{*}$ is equal to the codimension of $\mathcal{F}$. Hence, if $M$ has sectional curvature uniformly bounded below by $k\in\mathbb{R}$, and $\mathcal{F}$ is closed, then $M^{*}$ is an Alexandrov space with $\operatorname{curv}\geqslant k$.
2.3. Submetries from proper metric spaces
@@
Let $(X,d)$ be a metric space and let =$\mathcal{F}=\{L_{\alpha}\mid\alpha\in\Lambda\}$ be a partition of $X$. The foliation $\mathcal{F}$ is equidistant if, for any two distinct $L_{\alpha},L_{\beta}\in\mathcal{F}$, there exist $x_{\alpha}\in L_{\alpha}$ and $x_{\beta}\in L_{\beta}$ such that $d(x_{\alpha},x_{\beta})=d(L_{\alpha},L_{\beta})$. We will say that $\mathcal{F}$ is a metric foliation if it is an equidistant partition.
@@
When $X$ is a proper metric space, the leaves of a metric foliation of $X$ are closed subsets of $X$. Moreover, if $p\colon X\to Y$ is a submetry between metric spaces, then the partition =$\mathcal{F}=\{p^{-1}(y)\mid y\in Y\}$ is a metric foliation of $X$ with closed leaves (see [29, p. 19]).
Lemma 2.2 ([29, Lemma 4.7]).

### v3

@@ 2.1 Singular Riemannian foliations
    For each $p\in M$, there exists local smooth vector fields spanning the tangent spaces of the leaves.
@@
We call any leaf of maximal dimension a regular leaf; leaves that are not regular are called singular. Given any Riemannian manifold $M$, the foliation consisting of one leaf $\mathcal{F}=\{M\}$ and the foliation where each leaf consists of just a point $\mathcal{F}=\{\{x\}\mid x\in M\}$ are trivial examples of singular Riemannian foliations. Other examples of singular Riemannian foliations are given by the partition of a complete Riemannian manifold $M$ by the orbits of an isometric action of a compact Lie group. The partition of $M$ into the fibers of a Riemannian submersion $f\colon M\to N$} yields a further example of a singular Riemannian foliation. Note that there are infinitely many examples of singular Riemannian foliations which are not given by group actions nor from Riemannian submersions [21, 48].
@@
Let $(M,\mathcal{F})$ be a singular Riemannian foliation. Then we have a singular distribution $H\subset M$, called the horizontal distribution, given by setting $H_{x}=\nu_{x}L_{x}$, the normal tangent space to the leaf $L_{x}$ at $x$. The codimension of the foliation, denoted by $\operatorname{codim}(\mathcal{F})$, is the codimension of any regular leaf in $M$. We say $(M,\mathcal{F})$ is closed if all leaves are closed in $M$. The leaf space of the foliation is the set of equivalence classes $M^{\ast}=M/\mathcal{F}$ equipped with the quotient topology. We have a natural projection map $\pi\colon M\to M^{\ast}$ which is continuous with respect to the quotient topology. Given a subset $A\subset M$, we let $A^{\ast}=\pi(A)$.
@@
Let $(M,\mathcal{F})$ be a closed singular Riemannian foliation. Fix $x\in M$ and consider the normal space $\nu_{x}L_{x}$ to the tangent space $T_{x}L_{x}\subset T_{x}M$ at $x$. Next, for $\varepsilon>0$ sufficiently small, define $\nu_{x}^{\varepsilon}L_{x}=(\nu_{x}L_{x})\cap B_{\varepsilon}(0)$}, where $B_{\varepsilon}(0)$ is the closed ball of radius $\varepsilon$ in $T_{x}M$. Set $S_{x}=\exp_{x}(\nu_{x}^{\varepsilon}L_{x})$. The intersection of the leaves of $\mathcal{F}$ with $S_{x}$ induces a foliation $\mathcal{F}|_{S_{x}}$ on $S_{x}$ whose leaves are the connected components of the intersection between the leaves of $\mathcal{F}$ and $S_{x}$. Although $\mathcal{F}|_{S_{x}}$ may not be a singular Riemannian foliation with respect to the induced metric of $M$ on $S_{x}$ (the leaves of $\mathcal{F}|_{S_{x}}$ may not be equidistant with respect to the induced metric), the pull-back foliation $\mathcal{F}^{x}=\exp_{x}^{\ast}(\mathcal{F}|_{S_{x}})$} is a singular Riemannian foliation on $\nu_{x}^{\varepsilon}L_{x}$ equipped with the Euclidean metric $g^{\perp}_{x}:=(g_{x})|_{\nu_{x}L_{x}}$ (see [45, Proposition 6.5]). The foliation $(\nu_{x}^{\varepsilon}L_{x},\mathcal{F}^{x})$ is called the infinitesimal foliation at $x$.
@@
The infinitesimal foliation $(\nu_{x}^{\varepsilon}L_{x},\mathcal{F}^{x})$ is invariant under homotheties fixing the origin (see [45, Lemma 6.2]). Furthermore, the origin $\{0\}\subset\nu_{x}^{\varepsilon}L_{x}$ is a leaf of the infinitesimal foliation. Since the leaves of $\mathcal{F}^{x}$ are equidistant, the origin being a leaf implies that any leaf of $\mathcal{F}^{x}$ is at a constant distance from $\{0\}$. Therefore, each leaf of the infinitesimal foliation is contained in a round sphere centered at the origin. Hence, we may consider the infinitesimal foliation restricted to the unit normal sphere of $\nu_{x}L_{x}$, denoted by $S^{\bot}_{x}$, resulting in a foliated round sphere $\bigl(S^{\bot}_{x},\mathcal{F}^{x}\bigr)$ with respect to the standard round metric of $S^{\bot}_{x}$. This foliation is also called the infinitesimal foliation. Henceforth, when referring to the “infinitesimal foliation”, we will mean $\bigl(S^{\bot}_{x},\mathcal{F}^{x}\bigr)$. Note that there is no loss of generality in doing so, since $(\nu_{x}L_{x},\mathcal{F}^{x})$ is invariant under homothetic transformations and thus one may recover it from $\bigl(S^{\bot}_{x},\mathcal{F}^{x}\bigr)$.
Let $L$ be a closed leaf of a singular Riemannian foliation $(M,\mathcal{F})$, and $\gamma\colon[0,1]\to L$ a piecewise smooth curve with $\gamma(0)=x$. By [43], there exists a continuous map $G\colon[0,1]\times\nu_{p}L\to\nu L$ such that
  1. (a)
    $G(t,v)\in\nu_{\gamma(t)}L$ for every $(t,v)\in[0,1]\times\nu_{x}L$.
  2. (b)
    For every $t\in[0,1]$, the restriction $G|_{\{t\}\times\nu_{x}L}\colon\ \nu_{x}L\to\nu_{\gamma(t)}L$ is a linear isometry preserving the leaves of $\nu L$.
  3. (c)
    For every $s\in\mathbb{R}$, $\exp_{\gamma(t)}(sG(t,v))$ belongs to the same leaf as $\exp_{x}(sv)$.
@@
We denote by $\mathrm{O}\bigl(S^{\perp}_{x},\mathcal{F}^{x}\bigr)$ the group of foliated isometries of the infinitesimal foliation, i.e., the isometries which preserve the infinitesimal foliation. For each loop $\gamma$ at $x$, the map $G_{\gamma}\colon{\nu_{x}L\to\nu_{x}L}$ given by $G_{\gamma}(v)=G(1,v)$ is a foliated linear isometry (see [44, Corollary 15]). Therefore, we have a group homomorphism $\rho\colon\Omega(L,x)\to\mathrm{O}\bigl(S^{\perp}_{x},\mathcal{F}^{x}\bigr)$ from the loop space of $L_{x}$ at $x$ to the foliated isometries of the infinitesimal foliation by setting $\rho(\gamma)=G_{\gamma}$.
@@
An isometry in $\mathrm{O}\bigl(S^{\perp}_{x},\mathcal{F}^{x}\bigr)$ may map a leaf to a different leaf. By $\mathrm{O}(\mathcal{F}^{x})$ we denote the foliated isometries preserving the foliation, i.e., isometries $f\in\mathrm{O}\bigl(S^{\perp}_{x},\mathcal{F}^{x}\bigr)$ such that for any leaf $\mathcal{L}$ of $\bigl(S^{\perp}_{x},\mathcal{F}^{x}\bigr)$, we have $f(\mathcal{L})\subset\mathcal{L}$. The natural action of $\mathrm{O}\bigl(S^{\perp}_{x},\mathcal{F}^{x}\bigr)$ on the quotient $S^{\perp}_{x}/\mathcal{F}^{x}$ has kernel $\mathrm{O}(\mathcal{F}^{x})$. Using the fact that if two loops $\gamma_{1}$ and $\gamma_{2}$ based at $x$ are homotopic, then $G_{\gamma_{1}}^{-1}\circ G_{\gamma_{2}}$ is in the kernel of the action of $\mathrm{O}\bigl(S^{\perp}_{x},\mathcal{F}^{x}\bigr)$ on $S^{\perp}_{x}/\mathcal{F}^{x}$, one may show that there is a well-defined group homomorphism
$\rho\colon\ \pi_{1}(L,x)\to\mathrm{O}\bigl(S^{\perp}_{x},\mathcal{F}^{x}\bigr)/\mathrm{O}(\mathcal{F}^{x}),$
@@
given by $\rho[\gamma]=[G_{\gamma}]$ (see, for example, [13, Lemma 2.4 and Proposition 2.5]). We define the holonomy of the leaf $L$ as the image $\Gamma_{L}<\mathrm{O}\bigl(S^{\perp}_{x},\mathcal{F}^{x}\bigr)/\mathrm{O}(\mathcal{F}^{x})$ of $\pi_{1}(L,x)$ under the homomorphism $\rho$. When we consider the holonomy of a leaf $L_{x}$ through a point $x\in M$, we denote it by $\Gamma_{x}$. A regular leaf $L$ is a principal leaf if its holonomy group is trivial, and exceptional otherwise. The set $M_{\mathrm{prin}}\subset M$ consisting of the union of principal leaves is an open and dense subset of $M$ (see for example [13, Proposition 2.8]).
2.2 Alexandrov spaces
[…]
@@ 2.2 Alexandrov spaces
$\angle(c_{1},c_{2})=\lim_{s,t\to 0}\measuredangle(c_{1}(s),x,c_{2}(t)),$
provided the limit exists. When $X$ is an Alexandrov space and the curves $c_{1}$ and $c_{2}$ are geodesics, the angle $\angle(c_{1},c_{2})$ exists (see [11, Proposition 4.3.2]).
Given two geodesics $c_{1}\colon[0,1]\to X$ and $c_{2}\colon[0,1]\to X$ in an Alexandrov space $X$ with common start point $x\in X$, we say that $c_{1}$ is equivalent to $c_{2}$ if $\angle(c_{1},c_{2})=0$. Let $\widetilde{\Sigma}_{x}(X)$ the set of equivalence classes of geodesics starting at $x$. We define a metric on this set by setting the distance between two classes to be the angle formed between any two representatives of each class. The space of directions $\Sigma_{x}(X)$ of $X$ at $x$ is the metric completion of $\widetilde{\Sigma}_{x}(X)$. By [12, Corollaries 7.10 and 7.11], the metric space $\Sigma_{x}(X)$ is an Alexandrov space of $\operatorname{curv}\geq 1$.
By [11, Theorem 10.4.1], any Alexandrov space $\Sigma$ with $\operatorname{curv}\geq 1$ has $\operatorname{diam}(\Sigma)\leq\pi$. Let $(X,d)$ be a metric space with $\operatorname{diam}(X)\leq\pi$. The Euclidean cone over $X$, denoted by $CX$, is the set $X\times[0,\infty)/(x,0)\sim(y,0)$} equipped with the metric
@@
Given an Alexandrov space $X$ and a point $x\in X$, we define the tangent cone of $X$ at $x$ as $T_{x}X=C\Sigma_{x}(X)$. We denote the vertex of $T_{x}X$ by 0.
@@
The leaf space $M^{\ast}$ of a singular Riemannian foliation $(M,\mathcal{F})$ on a complete Riemannian manifold with closed leaves inherits a complete metric $d^{\ast}$ from $M$, known as the leaf-projection metric. For $x^{\ast},y^{\ast}\in M^{\ast}$, the distance $d^{\ast}(x^{\ast},y^{\ast})$ is defined as $d^{\ast}(x^{\ast},y^{\ast})=d(L_{x},L_{y})$, the distance between the leaves $L_{x}$ and $L_{y}$ considered as subsets of $M$. Equipped with the metric $d^{\ast}$, the leaf space $M^{\ast}$ has curvature locally bounded below in the triangle comparison sense discussed above. Specifically, if $U\subset M$ is an open neighborhood with sectional curvature bounded below by $k_{U}\in\mathbb{R}$, then the projection $U^{\ast}\subset M^{\ast}$ has curvature bounded below by $k_{U}$ (see [38]). The Hausdorff dimension of $M^{*}$ is equal to the codimension of $\mathcal{F}$. Hence, if $M$ has sectional curvature uniformly bounded below by $k\in\mathbb{R}$, and $\mathcal{F}$ is closed, then $M^{*}$ is an Alexandrov space with $\operatorname{curv}\geq k$.
2.3 Submetries from proper metric spaces
@@
Let $(X,d)$ be a metric space and $\mathcal{F}=\{L_{\alpha}\mid\alpha\in\Lambda\}$ a partition of $X$.
The foliation $\mathcal{F}$ is equidistant if for all leaves $L_{\alpha},L_{\beta}\in\mathcal{F}$ and all $x_{\alpha}\in L_{\alpha}$, we have $d(L_{\alpha},L_{\beta})=d(x_{\alpha},L_{\beta})$, where the first distance is the distance between subsets of $M$. We will say that $\mathcal{F}$ is a metric foliation if it is an equidistant partition.
@@
When $X$ is a proper metric space, the leaves of a metric foliation of $X$ are closed subsets of $X$. Moreover, if $p\colon X\to Y$ is a submetry between metric spaces, then the partition $\mathcal{F}=\big\{p^{-1}(y)\mid y\in Y\big\}$ is a metric foliation of $X$ with closed leaves (see [36, p. 19]).
Lemma 2.2 ([36, Lemma 4.7]).